# Multi-Objective Planning with Contextual Lexicographic Reward Preferences

### v1

@@ 4.2. Conflict Detection
6:$S_{solved}.\textsc{Append}(s_{g})$triangleright only $s_{g}$ is solved initially
7:$S_{unsolved}.\textsc{Remove}(s_{g})$
8:while $S_{unsolved}\not=\phi$ $\triangleright$ continue if unsolved states remain
@@
20:         if $\Delta(s)<\epsilon$ $\triangleright$ label $s$ as solved if $V_{r}(s)$ converged
@@
25:$S_{\emph{conflict}}\leftarrow[s\,|\,s\in S\text{ if }V_{r}(s)=0]$$\triangleright$ list of conflict states
@@
1:Input policy $\pi_{G}$, set of contexts $\mathcal{C}$, state to context mapping function $\mathcal{Z}$, context to objective ordering map $f_{\mathbf{w}}$, meta-ordering over contexts $\Omega\equiv c_{1}\succ\dots\succ c_{m}$, $S$, $A$, $T$, R, $\gamma$, goal $s_{g}$, list conflict states from ConflictChecker $S_{\emph{conflict}}$
@@
10:         $C_{update}\leftarrow C_{update}\cup c_{j}$triangleright contexts to be updated
@@ 5. Learning State-Context Mapping
where $Pr(\@vec{r}|s,a,c)$ is the probability of observing the reward vector ${r}$ when executing action $a$ in state $s$ under context $c$, $Pr(a|s,c,\boldsymbol{\tau})$ is the probability of the expert executing action $a$ in state $s$ if $c$ is the underlying context associated with $s$, and $Pr(c|\boldsymbol{\tau})$ denotes the prior probability, given $\boldsymbol{\tau}$. We marginalize over actions to ensure that the computed posterior reflects the overall probability of context $c$ across all potential actions that the expert could have taken, given the state and reward (which varies based on the context). We first describe how these terms are calculated for state-action pairs in the dataset and then discuss how to estimate the posterior for other states. Algorithm 3 outlines the steps involved in inferring $\mathcal{Z}$.
@@
The probability of observing a particular reward vector ${r}$ in the data is either one or zero, based on $f_{R}$ in the CLMDP (line 10). That is, $Pr(\@vec{r}|s,a,c)\!=\!1$ when ${r}$ matches the reward vector under context $c$ ($f_{R}(c)$), and $Pr(\@vec{r}|s,a,c)\!=\!0$ otherwise.
@@
1:Input $\mathcal{C}$, $\mathbf{o}$, $f_{\mathbf{w}}$, $\Omega$, $S$, $A$, $T$, $\mathbf{R}$, $s_{g}$, Set of expert trajectories $\boldsymbol{\tau}$
2:$\mathcal{P}(s,a)\leftarrow[],\forall(s,a)\in\boldsymbol{\tau}$
@@
4:     $\pi_{c}\leftarrow LVI(S,A,T,\mathbf{R},f_{\mathbf{w}}(c))$triangleright compute policy for $c$
@@
18:         Compute posterior $Pr(c|s,\@vec{r})$ using Eqn. 2
@@
For states not in $\boldsymbol{\tau}$, the posterior is estimated using $Pr(c|s,\@vec{r})\propto\sum\limits_{a\in A}Pr(\@vec{r}|s,a,c)\cdot Pr(a|s,c)\cdot Pr(c)$, where each term is calculated based on the policy computed for each context, instead of using expert data. The action probability is determined by the policy $\pi_{c}$ that is computed by considering all states mapped to context $c$. $Pr(a|s,c)\!=\!1$ if $\pi_{c}(s)=a$, and $Pr(a|s,c)\!=\!0$ otherwise (line 16). The prior $Pr(c)$ is a uniform distribution over all contexts (line 17). $Pr(\@vec{r}|s,a,c)$ is determined based on $f_{R}$, similar to estimating it under expert data (line 10).
@@ 6. Experiments
We evaluate our approach against six baselines using three domains in simulation and validate our results through hardware experiments. Unless specified otherwise, all algorithms were implemented by us in Python and the simulations were run on an iOS machine with $18$ GB RAM.  [^1] Code: https://tinyurl.com/Contextual-LMDP. [^2] Videos: https://tinyurl.com/CLMDP-Project-Page.
@@
Table 1. Percentage conflicts and task completion for multi-context techniques, averaged over $100$ trials in five instances.

### v2

@@ 4.2. Conflict Detection
[…]
@@
[…]
@@
[…]
@@
1:Input policy $\pi_{G}$, set of contexts $\mathcal{C}$, state to context mapping function $\mathcal{Z}$, context to objective ordering map $f_{\mathbf{w}}$, meta-ordering over contexts $\Omega\equiv c_{1}\succ\dots\succ c_{m}$, $S$, AA, $T$, R, $\gamma$, goal $s_{g}$, list conflict states from ConflictChecker $S_{\emph{conflict}}$
@@
10:   $C_{update}\leftarrow C_{update}\cup c_{j}$ $\triangleright$ contexts to be updated
@@ 5. Learning State-Context Mapping
where Pr(r|s,a,c)Pr(\@vec{r}|s,a,c) is the probability of observing the reward vector r\@vec{r} when executing action $a$ in state $s$ under context $c$, $Pr(a|s,c,\boldsymbol{\tau})$ is the probability of the expert executing action $a$ in state $s$ if $c$ is the underlying context associated with $s$, and $Pr(c|\boldsymbol{\tau})$ denotes the prior probability, given $\boldsymbol{\tau}$. We marginalize over actions to ensure that the computed posterior reflects the overall probability of context $c$ across all potential actions that the expert could have taken, given the state and reward (which varies based on the context). We first describe how these terms are calculated for state-action pairs in the dataset and then discuss how to estimate the posterior for other states. Algorithm 3 outlines the steps involved in inferring $\mathcal{Z}$.
@@
The probability of observing a particular reward vector r\@vec{r} in the data is either one or zero, based on $f_{R}$ in the CLMDP (line 10). That is, Pr(r|s,a,c)=1Pr(\@vec{r}|s,a,c)\!=\!1 when r\@vec{r} matches the reward vector under context $c$ ($f_{R}(c)$), and Pr(r|s,a,c)=0Pr(\@vec{r}|s,a,c)\!=\!0 otherwise.
@@
1:Input $\mathcal{C}$, $\mathbf{o}$, $f_{\mathbf{w}}$, $\Omega$, $S$, AA, $T$, $\mathbf{R}$, $s_{g}$, Set of expert trajectories $\boldsymbol{\tau}$
[…]
@@
4:  $\pi_{c}\leftarrow LVI(S,A,T,\mathbf{R},f_{\mathbf{w}}(c))$ $\triangleright$ compute policy for $c$
@@
18:   Compute posterior Pr(c|s,r)Pr(c|s,\@vec{r}) using Eqn. 2
@@
For states not in $\boldsymbol{\tau}$, the posterior is estimated using $Pr(c|s,\@vec{r})\propto\sum\limits_{a\in A}Pr(\@vec{r}|s,a,c)\cdot Pr(a|s,c)\cdot Pr(c)$, where each term is calculated based on the policy computed for each context, instead of using expert data. The action probability is determined by the policy $\pi_{c}$ that is computed by considering all states mapped to context $c$. Pr(a|s,c)=1Pr(a|s,c)\!=\!1 if $\pi_{c}(s)=a$, and Pr(a|s,c)=0Pr(a|s,c)\!=\!0 otherwise (line 16). The prior $Pr(c)$ is a uniform distribution over all contexts (line 17). Pr(r|s,a,c)Pr(\@vec{r}|s,a,c) is determined based on $f_{R}$, similar to estimating it under expert data (line 10).
@@ 6. Experiments
We evaluate our approach against six baselines using three domains in simulation and validate our results through hardware experiments. Unless specified otherwise, all algorithms were implemented by us in Python and the simulations were run on an iOS machine with $18$ GB RAM.  [^1] Code: https://tinyurl.com/Contextual-LMDP.
@@
Table 1. Percentage conflicts and task completion for multi-context techniques, averaged over 100100 trials in five instances.